# Theorem of Cassels and sums of squares of regular functions

### v1

@@ Abstract.
Let $C$ be an irreducible nonsingular curve over a formally real field $K$ and let $f$ be a regular function which is a sum of squares in the ring of regular functions of the surface $C\times\mathbb{A}_{K}$. We show that the (sum of squares) length of $f$ is the same in the ring of regular functions of $C\times\mathbb{A}_{K}$ and in the field of rational functions of $C\times\mathbb{A}_{K}$. With this we are able to show that the Pythagoras number of the ring of regular functions is the same as the Pythagoras number of the field of rational functions, for some product algebraic surfaces. This shows that the Pythagoras number of the ring of regular functions of the real plane is 4, which was not known before.
@@ Introduction
Let $f$ be a real valued function defined on a space $X$. Assume that $f$ is a sum of squares of functions on $X$. What is the minimal number of squares needed to represent $f$? In this paper we tackle this problem for regular functions on an algebraic surface, defined over a formally real field $K$. This problem is closely related to that of computing the Pythagoras number of a ring.
@@
For a commutative ring $R$ with identity, determining its Pythagoras number is a rather difficult task. As we will be considering real rings (i.e. rings in which $-1$ is not a sum of squares), let us discuss what is known in this case. It is known that any 1–dimensional affine $\mathbb{R}$–algebra has finite Pythagoras number (see [7]). On the other hand, any affine $K$–algebra over a formally real field $K$ of dimension at least 3 has infinite Pythagoras number. Similarly any regular local ring $R$ of dimension at least 3 with formally real residue field has $p(R)=\infty$ (see [7] for both results).
The case of dimension 2 seems to be the most mysterious one. It is known that $p(\mathbb{R}[[x,y]])=2$ and $p(\mathbb{Z}[[x]])=5$, but $p(\mathbb{Z}[x])=p(\mathbb{R}[x,y])=\infty$ (see [7]). Recently, there has been some development in this area. Pythagoras number of rings of the type $\mathbb{R}[x,y,\sqrt{f(x,y)}]$ where $f(x,y)$ is a polynomial satisfying some mild conditions has been computed in [4]. See also [1] for the $k$-regulous case and [27] for some other problems related to sums of squares. For a 2–dimensional regular local ring $R$ it was shown in [24] that finiteness of the Pythagoras number of the field of fractions of $R$ implies finiteness of the Pythagoras number of $R$ and explicit bound was given.
Computing the Pythagoras number of a number field is rather simple task [20, Chapter 7, Example 1.4]. On the other hand, computing the Pythagoras number of the ring of integers or orders is much more involved, see [9, 11, 16, 17, 18, 26] for some recent developments in this area. It is also worth to mention very recent paper [21], which deals with the Pythagoras number of the field of rational functions of a curve over a number field.
@@
Our main result relates the length of a regular function to the length of this function considered as a rational function. Surprisingly, the Pythagoras number was not known for a ring of regular functions of any surface, not even the real plane. Let us discuss what is known. If $R$ is a ”ring of regular functions” (RFR) in the sense of [19], then any totally positive unit can be written as a sum of squares, and explicit upper bound was given in terms of transcendence degree. For an algebraic surface over a formally real field $K$, the usual ring of regular functions is an RFR, hence the result of Mahe applies. This is however far from computing the Pythagoras number, as there is no control on the lengths of non–units. In a similar spirit one defines the unital Pythagoras number i.e. one considers only units which are sums of squares and not general elements which are sums of squares. It is worth to mention the paper [3, Theorem 5.17] where this notion was studied. The most interesting part is the relation with higher unital Pythagoras numbers. Such relation is known only in very specific cases for usual Pythagoras numbers [15, Conjecture 5.2].
Let us describe the content of the paper. In Section 1 we recall necessary facts from algebraic geometry and theory of quadratic forms needed to prove the main theorem. In Section 2 we provide applications of main result to the problem of Pythagoras numbers. We compute the Pythagoras number of the ring of regular functions on a real plane, which turns out to be 4.
1. Main result
Let us start with preliminary definitions.
@@ 1. Main result
Let $K$ be a formally real field. We will recall all the definitions regarding algebraic geometry and rings of regular functions. In this paper we will avoid the scheme theoretic point of view on algebraic varieties. For the treatment of algebraic sets over a real closed field see [5, Chapter 3]. Let $X\subset K^{n}$ be an algebraic set.
@@
where $\mathcal{I}(X)$ is the ideal of all the polynomials vanishing on $X$.
If $X$ is an irreducible algebraic set, then we denote its field of fractions by $K(X)$.
@@
We say that $x_{0}\in X$ is a nonsingular point if the local ring $K[X]_{\mathfrak{m}_{x_{0}}}$ is a regular local ring.
We say that an algebraic set $X\subset K^{n}$ is nonsingular if every point point is nonsingular.
As we will be interested only in $K$-rational points, we may allow singularities to occur in the algebraic closure of $K$. The crucial object of our study is the following
@@
Let us note that similarly as in the scheme theory one can consider sheaves of modules over the sheaf of regular functions. This geometric approach differs substantially from the case of an algebraically closed field (or scheme theoretic point of view). For example, over $\mathbb{R}$, Cartan’s Theorems $A$ and $B$ do not hold, however they do hold after a suitable multiblowup [12, 13].
In algebraic geometry one usually defines the notion of a regular function locally. However, since we deal only with formally real fields and by [5, Proposition 3.2.3] regular functions on $X$ are precisely those that admit a global denominator. In particular, $\mathcal{O}(X)$ is a localization of $K[X]$ at a set consisting of polynomial functions vanishing nowhere on $X$. If $K=\mathbb{R}$ and $X$ is a smooth algebraic set, then we localize at all polynomial functions which does not change sign on $X$. For more general field this is not true, as the following example shows.
@@
Let $K=\mathbb{Q}$ and $X=\mathbb{Q}$. Then the function $\frac{1}{x^{2}-2}$ belongs to $\mathcal{O}(X)$ even though it changes sign on $X$.
We may now define the notion of a length, which will be useful in the study of sums of squares.
Definition 1.6.
Let $R$ be a commutative ring with identity. We define the length of an element $a\in R$ to be the smallest positive integer $g$ such that $a$ can be written as a sum of at most $g$ squares.
@@
Let $R=\mathbb{Z}$ and $a=7$. In this case it is known that the length of $7$ is $4$.
From the very definition of the length we obtain the following fact
@@
Apart from algebraic geometry, we will use elements of the theory of quadratic forms (see [20]). Let us now recall the Theorem of Cassels [6], where only sums of squares were considered instead of a general quadratic form. The statement we present below can be found in [20, Chapter 1, Theorem 2.2].
@@
Let $K$ be a formally real field. We will consider product surfaces, hence to avoid awkward formulas such as $K[C\times K]$, we will denote the affine line over $K$ as $\mathbb{A}_{K}$. In this case, the above coordinate ring will be denoted as $K[C\times\mathbb{A}_{K}]$ which turns out to be $K[C][y]$, where $y$ is a single variable.
Lemma 1.10.
Let $C$ be a curve over $K$ and let $x_{0}\in C$ be a nonsingular point. Then the curve $D=\{x_{0}\}\times\mathbb{A}_{K}\subset C\times\mathbb{A}_{K}$ is a zero set of a single polynomial functions. Moreover, this polynomial function can be chosen to be the generator of the prime ideal of all polynomial functions vanishing on $D$.
@@
The coordinate ring of the surface $C\times\mathbb{A}_{K}$ is equal to $K[C][y]$ for a single variable $y$. One can find a function $g\in K[C]$ such that $g$ generates the maximal ideal $\mathfrak{m}_{x_{0}}$ of the point $x_{0}$. Then the function $g$ vanishes precisely on $\{x_{0}\}\times\mathbb{A}_{K}$. Additionally, $g$ is the generator of the prime ideal of polynomial functions vanishing on $D$.
We are now ready to prove the main result.
@@
Let $K$ be a formally real field. Let $C$ be an irreducible nonsingular curve over $K$, and $X=C\times\mathbb{A}_{K}$ be a product surface. Let $f\in\mathcal{O}(X)$ be a regular function which is a sum of squares. Then the lengths of $f$ in $K(X)$ and $\mathcal{O}(X)$ are equal.
@@
Let $f\in\mathcal{O}(X)$ be a sum of squares. By Lemma 1.8 the length of $f$ in $\mathcal{O}(X)$ is greater than or equal to the length of $f$ in $K(X)$.
By the previous lemma, the curve of the form $\{x_{0}\}\times\mathbb{A}_{K}$ is cut out globally by a single equation for any $x_{0}\in C$. We may write $f=f_{1}h^{2}$, where the zero set of $h$ consists precisely of finite number of curves of the aforementioned type. Replacing $f$ by $f_{1}$ we may assume that $f$ does not not vanish on any curve of the form $\{x_{0}\}\times\mathbb{A}_{K}$ for any $x_{0}\in C$.
Of course, $K[C][y]\subset K(C)[y]$. Denote by $N$ the length of $f$ in $K(X)$ and let $\varphi=\sum_{i=1}^{N}x_{i}^{2}$ be a quadratic form. By the very definition of $N$, $\varphi$ represents $f$ over $K(X)$, hence, by the Theorem of Cassels also over $K(C)[y]$. Let $f=\sum_{i=1}^{N}\frac{f_{i}^{2}}{g^{2}}$ be a presentation such that $f_{i}\in K[X]$ and $g\in K[C]$. Any zero $x_{0}\in C$ of $g$ corresponds to a curve of the form $\{x_{0}\}\times\mathbb{A}_{K}$ in $C\times\mathbb{A}_{K}$ and those are given by a single equation by Lemma 1.10. Hence, we can factor them from $g$ and all $f_{i}$ for $i=1,2,\dots,N$. Consequently, $g$ does not vanish and so we obtain representation of $f$ as a sum of squares of regular functions of length $N$. ∎
2. Applications and further remarks.
In this section we will apply the above result to the theory of Pythagoras numbers. We will also discuss possible generalization of the main result to the case of more general rings as well as to higher even powers instead of squares.
2.1. Pythagoras numbers
Corollary 2.1.
With the assumptions of Theorem 1.11 we obtain $p(\mathcal{O}(X))=p(K(X))$.
Proof.
It is enough to note that if $f\in K(X)$ is a rational function of length $N$, then one has a presentation $f=\frac{\sum_{i=1}^{N}f_{i}^{2}}{g^{2}}$ for $f_{i},g\in K[X]$. Necessarily, the lengths of $f$ and $fg^{2}$ are equal in $K(X)$ and additionally $fg^{2}\in K[X]$. By Theorem 1.11, the lengths of $fg^{2}$ in $\mathcal{O}(X)$ and $K(X)$ are equal, hence the thesis.
The above corollary reduces the problem of finding the Pythagoras number of the ring of regular functions to that of the field of rational functions. In some very specific cases this is known. We are able to deduce the following.
Corollary 2.2.
For $X=\mathbb{R}^{2}$ we have $p(\mathcal{O}(\mathbb{R}^{2}))=4$.
Results of Scheiderer [25] imply the following.
Theorem 2.3.
Let $p\in\mathbb{R}[x,y]$ be a nonnegative polynomial. Then, for sufficiently large $N$ the polynomial $(x^{2}+y^{2}+1)^{N}p$ is a sum of squares of polynomials.
However, no bound on the number of squares is known in this case. By the proof of Theorem 1.11 and Corollary 2.2 we able to refine this result.
Corollary 2.4.
Let $p\in\mathbb{R}[x,y]$ be a nonnegative polynomial. Then, there exists a polynomial $h\in\mathbb{R}[x]$ such that $h^{2}p$ is a sum of 4 squares of polynomials.
We get the precise number of summands at the cost of not knowing the precise denominator, which depends only on 1 variable.

### v2

# Sums of squares of regular functions on rational surfaces

@@ Abstract.
We study the sums of squares on cylinders of the form $X\times\mathbb{A}_{K}$ for a (weakly) factorial curve $C$. We prove the equality of the Pythagoras numbers of the ring of regular functions on the cylinder with that of the field of rational functions. We then apply these results to the case of (uniformly) rational varieties. We show that if $X$ is a nonsingular rational algebraic surface over the reals, then the Pythagoras number of the ring of regular functions on $X$ is bounded above by 12.
@@ Introduction
Let $f$ be a real valued function defined on a space $X$. Assume that $f$ is a sum of squares of functions on $X$. What is the minimal number of squares needed to represent $f$? In this paper we tackle this problem for regular functions on an algebraic surface defined over a formally real field $K$. This problem is closely related to that of computing the Pythagoras number of a ring.
@@
For a commutative ring $R$ with identity, determining its Pythagoras number is a rather difficult task. As we will be considering real rings (i.e. rings in which $-1$ is not a sum of squares), let us discuss what is known in this case. Any 1–dimensional affine $\mathbb{R}$–algebra has finite Pythagoras number (see [17]). On the other hand, any affine $K$–algebra over a formally real field $K$ of dimension at least 3 has infinite Pythagoras number. Similarly, any regular local ring $R$ of dimension at least 3 with formally real residue field has $p(R)=\infty$ (see [17] for both results). Let us also mention a very recent preprint [37] which deals with similar problems.
The case of dimension 2 seems to be the most mysterious. It is known that $p(\mathbb{R}[[x,y]])=2$ and $p(\mathbb{Z}[[x]])=5$, but $p(\mathbb{Z}[x])=p(\mathbb{R}[x,y])=\infty$ (see [17]). Recently, there has been some development in this area. Pythagoras number of rings of the type $\mathbb{R}[x,y,\sqrt{f(x,y)}]$, where $f(x,y)$ is a polynomial satisfying some mild conditions, has been computed in [10]. See also [2] for the $k$-regulous case, and [47] for some other problems related to sums of squares. For a 2–dimensional regular local ring $R$ it was shown in [43] that finiteness of the Pythagoras number of the field of fractions of $R$ implies finiteness of the Pythagoras number of $R$ and explicit bound was given. We are able to strengthen this result in some cases, see Corollary 3.7. See also [5] for the case of the field of fractions of formal power series over $\mathbb{R}$.
Computing the Pythagoras number of a number field is a rather simple task [38, Chapter 7, Example 1.4]. On the other hand, computing the Pythagoras number of the ring of integers or orders is much more involved, see [19, 20, 26, 31, 32, 33, 46] for some recent developments in this area. It is also worth to mention very recent paper [39], which deals with the Pythagoras number of the field of rational functions of a curve over a number field. This result has already been improved in [8].
@@
Our main result relates the length of a regular function to the length of this function considered as a rational function. Surprisingly, the Pythagoras number was not known for a ring of regular functions of any surface, not even the real plane. Let us discuss what is known. If $R$ is a ”ring of regular functions” (RFR) in the sense of [35], then any totally positive unit can be written as a sum of squares, and explicit upper bound was given in terms of transcendence degree. For an algebraic surface over a formally real field $K$, the usual ring of regular functions is an RFR, hence the result of Mahé applies. This is however far from computing the Pythagoras number, as there is no control on the lengths of non–units. In a similar spirit one defines the unital Pythagoras number i.e. one considers only units which are sums of squares and not general elements which are sums of squares. It is worth mentioning the paper [7, Theorem 5.17] where this notion was studied. The most interesting part is the relation with higher unital Pythagoras numbers. Such relation is known only in very specific cases for usual Pythagoras numbers [30, Conjecture 5.2].
Let us describe the content of this paper. In Section 1 we recall necessary facts from algebraic geometry. Section 2 contains discussion of (weakly) factorial varieties. We also study sums of squares in rings of regular functions on curves. In the next section we study sums of squares on cylinders i.e. products of a curve and a line. Section 4 is devoted to the study of (uniformly) rational surfaces. Among other results, we show that if $X$ is a smooth nonsingular rational surface over $\mathbb{R}$ then $p(\mathcal{O}(X))\leq 12$. We finish the paper with discussion of possible further generalizations of theorem of Cassels, and list some conjectures and open problems which may stimulate further research.
1. Preliminaries
Let us begin with some preliminary definitions. We will start with the notion of length, which will be useful in the study of sums of squares. Let $R$ be a commutative ring with identity.
Definition 1.1.
We define the length of an element $a\in R$ to be the smallest positive integer $g$ such that $a$ can be written as a sum of at most $g$ squares. If $a$ is not a sum of squares, then its length is infinite.
Example 1.2.
Let $R=\mathbb{Z}$ and $a=7$. In this case it is known that the length of $7$ is $4$.
From the very definition of length we obtain the following fact
Lemma 1.3.
This in particular implies that $p(R)\geq p(K)$ where $K$ is the fraction field of $R$.
@@ 1. Preliminaries
Let $K$ be a formally real field. We will recall all the definitions regarding algebraic geometry and rings of regular functions. In this paper we will avoid the scheme theoretic point of view on algebraic varieties. For the treatment of algebraic sets over a real closed field see [11, Chapter 3]. Let $X\subset K^{n}$ be a Zariski closed algebraic set.
@@
where $\mathcal{I}(X)$ is the ideal of all polynomials vanishing on $X$.
If $X$ is an irreducible algebraic set, then we denote field of fractions of $K[X]$ by $K(X)$.
@@
We say that $x_{0}\in X$ is a nonsingular point if the local ring $K[X]_{\mathfrak{m}_{x_{0}}}$ is a regular local ring. We say that an algebraic set $X\subset K^{n}$ is nonsingular if every point point is nonsingular.
As we will be interested only in $K$-rational points, we may allow singularities to occur in the algebraic closure of $K$. If $K$ is not algebraically closed field, then the coordinate ring lack many key properties. Usually, in real algebraic geometry, a much more natural ring is the ring of regular functions – the crucial object of our study.
@@
Let us note that similarly as in the scheme theory one can consider sheaves of modules over the sheaf of regular functions. Our geometric approach differs substantially from the case of an algebraically closed field (or scheme theoretic point of view). For example, for real algebraic sets with the sheaf of regular functions, Cartan’s Theorems $A$ and $B$ do not hold, however they do hold after a suitable multiblowup [27, 28].
In algebraic geometry one defines the notion of a regular function locally. However, since we deal only with formally real fields, by [11, Proposition 3.2.3] regular functions on $X$ are precisely those that admit a global denominator. In particular, $\mathcal{O}(X)$ is a localization of $K[X]$ at a set consisting of polynomial functions vanishing nowhere on $X$.
Remark 1.8.
By an algebraic variety we will mean an irreducible algebraic set.
Remark 1.9.
In this paper we will consider product surfaces, hence to avoid awkward formulas such as $K[C\times K]$, we will denote the affine line over $K$ as $\mathbb{A}_{K}$.
Let $X\subset K^{n}$ be an irreducible algebraic set.
Definition 1.10.
We say that $X$ is a factorial variety if the ring $K[X]$ is a unique factorization ring. We say that $X$ is a weakly factorial variety if the ring $\mathcal{O}(X)$ is a unique factorization ring.
If $X$ is a factorial variety then it is obviously weakly factorial. Also, by a classical result in algebra, if $X$ is factorial, the so is cylinder $X\times\mathbb{A}_{K}$ over $X$. We will now discuss properties of weakly factorial varieties.
Lemma 1.11.
Let $X$ be a weakly factorial variety. Then the cylinder $X\times\mathbb{A}_{K}$ is a weakly factorial variety.
Proof.
By the assumption $\mathcal{O}(X)$ is a UFD, hence so is $\mathcal{O}(X)[y]$ by [24, Theorem 6.14]. Of course $K[X\times\mathbb{A}_{K}]\subset\mathcal{O}(X)[y]$. Let $S$ be the multiplicative set of all polynomial functions vanishing nowhere on $X\times\mathbb{A}_{K}$. Straightforward calculation shows that $S^{-1}K[X\times\mathbb{A}_{K}]=\mathcal{O}(X\times\mathbb{A}_{K})=S^{-1}\mathcal{O}(X)[y]$. As a consequence, $\mathcal{O}(X\times\mathbb{A}_{K})$ is a UFD as a localization of a UFD. ∎
@@
It is known that 1-dimensional unit sphere $\mathbb{S}^{1}$ is not a factorial variety, while 2-dimensional unit sphere $\mathbb{S}^{2}$ is, both considered as algebraic sets over $\mathbb{R}$ (see [45, Theorem 5]). The hyperbola $X=\{xy=1\}\subset K^{2}$ is also factorial, since its ring of polynomial functions $K[X]\cong K[x,x^{-1}]$ is a localization of $K[x]$.
2. Weakly factorial varieties over $\mathbb{R}$.
If we are working over the field of real numbers, we are able to give equivalent conditions for weak factoriality. In particular, we have
Theorem 2.1.
[11, Proposition 12.4.14] Let $X$ be a nonsingular compact algebraic variety over $\mathbb{R}$ of dimension $d$. Then the following conditions are equivalent:
  • $X$ is weakly factorial,
  • $H_{d-1}^{alg}(X,\mathbb{Z}/2)=0$.
Here, $H_{d-1}^{alg}(X,\mathbb{Z}/2)$ is the group of $d-1$ homology classes generated by classes represented by compact algebraic subsets. The above Theorem holds, for example if $H_{d-1}(X,\mathbb{Z}/2)=0$. As pointed out by Shiota, the factoriality of $X$ is not a topological invariant as the following example shows.
@@
[44, Exemple] Consider two algebraic sets, $X=\{x^{2}+y^{2}+z^{2}=1\}$ and $Y=\{x^{4}+y^{4}+z^{2}=1\}$. $X$ and $Y$ are homeomorphic, both are weakly factorial and $X$ is factorial while $Y$ is not. To see this, note that in $\mathbb{R}[Y]$ we have $x^{4}+y^{4}=1-z^{2}=(1-z)(1+z)=(x^{2}+\sqrt{2}xy+y^{2})(x^{2}-\sqrt{2}xy+y^{2})$ (cf. [12, Section 7]).
Non compact case was also studied, however much less is known (cf. [12, Theorem 5, Remark 6]). In particular, if $X$ is a nonsingular nonorientable algebraic set, then $X$ is not weakly factorial.
In this paper, weakly factorial curves will be of great importance. Let us now provide a topological classification of those.
@@
[21, Theorem 2.3] Let $X$ be a nonsingular algebraic curve over $\mathbb{R}$. Assume that $X$ has a compact connected component. Then $X$ is not a weakly factorial variety, hence it is not factorial.
In its original statement, the above Theorem shows that the maximal ideal corresponding to a point in a compact connected component cannot be principal. This is shown for the coordinate ring $\mathbb{R}[X]$ as well as $S^{-1}\mathbb{R}[X]$ where $S$ is the multiplicative set of all polynomial functions vanishing nowhere on the given connected component. In particular, this ring is a localization of the ring of regular functions $\mathcal{O}(X)$, hence $X$ is not weakly factorial.
Theorem 2.4.
Let $X$ be a nonsingular algebraic curve over $\mathbb{R}$ with no compact connected component. Then $X$ is weakly factorial.
Proof.
First of all, let us note that a nonsingular real algebraic curve is homeomorphic to a disjoint union of circles and intervals. Since $X$ has no compact connected component, $X$ is homeomorphic to a disjoint sum of finitely many open interval, hence $H^{1}(X,\mathbb{Z}_{2})=0$. Secondly, let us embed $X\subset\mathbb{P}^{n}(\mathbb{R})$ for some $n$. Denote by $X^{\prime}$ the possible desingularization of the Zariski closure of the image of $X$ in $\mathbb{P}^{n}(\mathbb{R})$. In this case $X^{\prime}\setminus X$ consists of finitely many points. Moreover every class in the homology group of degree 00, i.e. $H_{0}(X^{\prime}\setminus X,\mathbb{Z}_{2})$ is represented by a compact algebraic subset. The result now follows readily from the theorem below.
Theorem 2.5.
[44, Théorème 2] Let $X$ be an irreducible, non compact algebraic set of dimension $n$ with $H^{1}(X,\mathbb{Z}_{2})=0$. If every element of $H_{n-1}(X^{\prime}\setminus X,\mathbb{Z}_{2})$ is realized by a compact algebraic set then $X$ is weakly factorial.
The above results can be combined into the following.
Theorem 2.6.
Let $X$ be a nonsingular algebraic curve over $\mathbb{R}$. Then $X$ is weakly factorial if and only if $X$ has no compact connected component.
We see that for a nonsingular irreducible curve $C$ over $\mathbb{R}$, weak factoriality is a local property in the sense that any point $x\in C$ has a Zariski open neighborhood $U$ such that $\mathcal{O}(U)$ is a weakly factorial curve.
Remark 2.7.
The results of this Section essentially use the properties of the real numbers. In would be interesting in its own right to find a generalization of the above to arbitrarily fields, cf. Section 6.
Let us a finish this section with a discussion of sums of squares of regular functions on curves. Let $K$ be a formally real field.
Proposition 2.8.
Let $C\subset K^{n}$ be an irreducible weakly factorial curve. Then $p(\mathcal{O}(C))=p(K(C))$.
Proof.
Since $K(C)$ is the field of fractions of $\mathcal{O}(C)$ it is enough to show that the length of a regular function $f$ is the same in both aforementioned rings. Let $f\in\mathcal{O}(C)$ be a sum of squares. By the weak factoriality, we may factor $f$ as $f=g_{1}^{2}g$, where $g_{1},g\in\mathcal{O}(C)$ and $g$ is a unit. Write now $g$ as a sum of squares in $K(C)$ i.e. $g=\frac{\sum_{i=1}^{N}f_{i}^{2}}{h^{2}}$ with $h,f_{i}\in K[C]$ for $i=1,2,\dots,N$. By the weak factoriality, any possible zero of $h$ could be factored out from all of the $f_{i}$’s. We then obtain a representation of $g$ with a possibly smaller number of summands, all of which are regular functions. Hence the length of $f$ in $\mathcal{O}(C)$ and $K(C)$ are equal.
We may strengthen the above result for $K=\mathbb{R}$.
Theorem 2.9.
If $C$ be a nonsingular irreducible algebraic curve over $\mathbb{R}$ then $p(\mathcal{O}(C))\leq 4$.
Proof.
The result of Pfister [38, Corollary 3.4] implies that $p(\mathbb{R}(C))\leq 2.$ By the Theorem 2.6 we may take a covering of $C$ by two open subsets $U_{1},U_{2}\subset C$ such that $\mathcal{O}(U_{i})$ is a UFD. Let $f\in\mathcal{O}(C)$ be a sum of squares. Choose functions $h_{1},h_{2}\in\mathcal{O}(C)$ such that $\mathcal{Z}(h_{i})=C\setminus U_{i}$. By the previous proposition, we have that $f_{i}=f|_{U_{i}}$ is expressible as a sum of at most two squares in $\mathcal{O}(U_{i})$, i.e. $f_{i}=\frac{f_{1i}^{2}+f_{i2}^{2}}{g_{i}^{2}}=\frac{(f_{1i}^{2}+f_{i2}^{2})h_{i}^{2}}{g_{i}^{2}h_{i}^{2}}$, where the zero set of $g_{i}$ is contained in $C\setminus U_{i}$ and $f_{ij}\in\mathcal{O}(U_{i})$, $i,j=1,2$. Having those two presentations we may add numerators and denominators together to obtain a global presentation of $f$ i.e.
$f=\frac{(f_{11}^{2}+f_{12}^{2})h_{1}^{2}+(f_{21}^{2}+f_{22}^{2})h_{2}^{2}}{g_{1}^{2}h_{1}^{2}+g_{2}^{2}h_{2}^{2}}=\frac{\left((f_{11}^{2}+f_{12}^{2})h_{1}^{2}+(f_{21}^{2}+f_{22}^{2})h_{2}^{2}\right)(g_{1}^{2}h_{1}^{2}+g_{2}^{2}h_{2}^{2})}{(g_{1}^{2}h_{1}^{2}+g_{2}^{2}h_{2}^{2})^{2}}.$
By construction, the denominator does not vanish on $C$. Applying the 2-square (or 4-square) identity in the numerator finishes the proof. ∎
In the above proof we have shown the existence of a presentation as a quotient of regular functions, however, we have defined regular functions as quotients of polynomial functions. Since the denominators of global regular functions cannot vanish, one can always simplify such presentations to the desired form if necessary.
Remark 2.10.
The above results can be also shown for higher even Pythagoras numbers (cf. [30]). The proofs are essentially the same, hence we will omit them.
3. Sums of squares on cylinders.
As we have discussed the case of curves, the main aim of this section is to generalize Proposition 2.8 to a possibly large class of algebraic surfaces.
@@
Apart from algebraic geometry, we will use elements of the theory of quadratic forms (see [38]). Let us now recall the Theorem of Cassels [16], where only sums of squares were considered instead of a general quadratic form. The statement we present below can be found in [38, Chapter 1, Theorem 2.2].
@@
We are now ready to prove the main result of this Section.
Theorem 3.2.
Let $K$ be a formally real field. Let $C$ be an irreducible nonsingular weakly factorial curve over $K$, and $X=C\times\mathbb{A}_{K}$ be a product surface. Then $p(\mathcal{O}(X))=p(K(X))$.
@@
Similarly as before, since $K(X)$ is the field of fractions of $\mathcal{O}(X)$, it is enough to show that the lengths of a regular functions $f$ in $K(X)$ and $\mathcal{O}(X)$ are equal. By the weak factoriality of $X$, we may assume that the zero set of $f$ is finite and $f$ does not vanish on any curve. Denote the length of $f$ in $K(X)$ by $N$.
Of course, $K[C][y]\subset K(C)[y]$ and let $\varphi=\sum_{i=1}^{N}x_{i}^{2}$ be a quadratic form. By the very definition of $N$, $\varphi$ represents $f$ over $K(X)$, hence, by the Theorem of Cassels also over $K(C)[y]$. Let $f=\sum_{i=1}^{N}\frac{f_{i}^{2}}{g^{2}}$ be a presentation such that $f_{i}\in K[X]$ and $g\in K[C]$. If $g$ does not vanish on $C$ then we are done. Any zero $x_{0}\in C$ of $g$ corresponds to a curve of the form $\{x_{0}\}\times\mathbb{A}_{K}$ in $X$ and those are given by a single equation in $\mathcal{O}(X)$ by the weak factoriality of $C$ and $X$. Hence, we can factor them from $g$ and all $f_{i}$ for $i=1,2,\dots,N$. However, since the factorization takes places in $\mathcal{O}(X)$, it might happen that after this procedure the denominator $g$ might be replaced by a nowhere vanishing function on the whole $X$, not necessarily in $K[C]$, but in any case, $g$ does not vanish and so we obtain representation of $f$ as a sum of squares of regular functions of length $N$.
Remark 3.3.
If we further assume that the curve $C$ is factorial, then in the above proof we may assume that the denominator $g$ does not depend on $y$. This follows from the fact, that we are able to cancel out factors in the ring $K[X]$ rather than in $\mathcal{O}(X)$.
Corollary 3.4.
With the assumptions as in Theorem 3.2. Let $U\subset X$ be a Zariski open subset of $X$. Then $p(\mathcal{O}(X))=p(\mathcal{O}(U))$.
Proof.
If $K$ is not an algebraically closed field then any Zariski open subset is isomorphic to an algebraic set. In this case $\mathcal{O}(U)$ is isomorphic with a localization of $\mathcal{O}(X)$ and so we can derive $p(\mathcal{O}(X))\geq p(\mathcal{O}(U))\geq p(K(X))$. By the previous theorem the proof is done. ∎
Two above results immediately imply the following
Corollary 3.5.
Let $U\subset\mathbb{R}^{2}$ be a Zariski open subset. Then $p(\mathcal{O}(U))=p(\mathcal{O}(\mathbb{R}^{2}))=4$.
Similarly as for curves, we may strengthen Theorem 3.2 for cylinders over $\mathbb{R}$.
@@
If $C$ is a nonsingular irreducible algebraic curve over $\mathbb{R}$, then ${p(\mathcal{O}(C\times\mathbb{R}))\leq 8}$.
@@
It is enough to combine Theorem 3.2 and the proof of Theorem 2.9 with the result of Pfister [38, Corollary 3.4] which implies that $p(\mathbb{R}(C\times\mathbb{R}))\leq 4.$